An object having a mass of 10 g and a charge of + 8.0x10^-5 C is placed in an electric field defined by Ex = 3.0x10^3 N/C, Ey = - 6000 N/C, and Ez = 0. (a) What are the magnitude and direction of the force on the object? (b) If the object starts from rest at the origin, what will be its coordinates after 3.0 s?


1
Expert's answer
2021-11-01T17:14:46-0400

Mass(m)=10g

Charge (wa)="8.0\times10^{-5}C"

Electric field


"E_x=3\times10^{3}N/C,E_y=-{6000}N/C,E_z=0N/C"

"E=E_x\hat{i}+E_y\hat{j}+E_z\hat{k}"

"E=3000\hat{i}-6000\hat{j}+0\hat{k}"

Force

"F=qE"


"F=8.0\times10^{-5}\times(3000\hat{i}-6000\hat{j}+0\hat{k})"

"F=(0.24\hat{i}-0.48\hat{j} )N"


"|F|=\sqrt{(0.24^2++(-0.48)^2}=\sqrt{0.288}=0.5366N"

Direction


"\theta=tan^{-1}(\frac{-0.48}{0.24})=tan^{-1}({-2})=-63.43°"

Part(b)

Force (F)=ma

"a=\frac{F}{m}"

"a=\frac{0.5366}{10^{-2}}=53.66m/sec"

We know that

"S=ut+\frac{1}{2}at^2"


"X=0+\frac{1}{2}\times53.66\times3^2=241.47 m"